# Enhancing Breast Cancer Diagnosis in Mammography: Evaluation and Integration of Convolutional Neural Networks and Explainable AI

### v2

@@ 6 Conclusion
The study has explored the symbiosis of Convolutional Neural Networks (CNNs) and Explainable Artificial Intelligence (XAI) within the context of mammography for breast cancer diagnosis. The core of our research did not only refine the predictive accuracy of CNNs through the deployment of a fine-tuned ResNet50 model but also highlighted the blackbox, a void often criticized in deep learning methodologies by integrating XAI techniques—namely Grad-CAM, LIME, and SHAP. These methods addressed the ”black-box” nature of deep learning, offering insightful visual and quantitative interpretations of the model’s decision-making processes.
@@
Despite the promising results, this study acknowledges the limitations inherent in our approach, including the variability in explanation consistency observed with certain XAI techniques, notably LIME. Such limitations highlight the necessity for continuous refinement of both deep learning models and explanation methodologies to ensure that they meet the rigorous demands of clinical application.
In future, we have envisioned several trajectories for future research. Firstly, exploring additional deep learning architectures and training techniques could further enhance diagnostic accuracy. Secondly, developing more robust and consistent XAI methodologies will be crucial in improving the reliability of AI explanations. Thirdly, collaboration with medical professionals in designing AI tools can ensure that these technologies align with clinical needs and practices. Finally, extending our research to include multi-modal data, such as incorporating patient history and genetic information, could offer a more holistic approach to breast cancer diagnosis.

### v3

@@ 6 Conclusion
The study has explored the symbiosis of CNNs and XAI within the context of mammography for breast cancer diagnosis. The core of our research did not only improved the predictive accuracy of CNNs through the deployment of a fine-tuned ResNet50 model but also highlighted the blackbox, a void often criticized in DL methodologies, by integrating XAI techniques—namely Grad-CAM, LIME, and SHAP. These methods addressed the ”black box” nature of DL, offering insightful visual and quantitative interpretations of the model’s decision-making processes.
@@
Despite the promising results, this study acknowledges the limitations inherent in our approach, including the variability in explanation consistency observed with certain XAI techniques, notably LIME. Such limitations highlight the necessity for continuous refinement of both models and explanation methodologies to ensure that they meet the rigorous demands of clinical application.
In the future, we have envisioned several trajectories for future research. Firstly, exploring more DL architectures and training techniques could further enhance diagnostic accuracy. Secondly, developing more robust and consistent XAI methodologies will be crucial to improving the reliability of AI explanations. Thirdly, collaboration with medical professionals in designing AI tools can ensure that these technologies align with clinical needs and practices. Finally, extending our research to include multi-modal data, such as patient history and genetic information, could offer a more holistic approach to breast cancer diagnosis.